OpenStudy (calculusxy):

MEDAL for help in Algebra 1

OpenStudy (calculusxy):

Please don't mind the gibberish math around the problem :)

Vocaloid (vocaloid):

rate of change formula from a to b is f(b) - f(a) divided by (b-a)

Vocaloid (vocaloid):

so, since we want the change from x = 0 to x = 3, we use the formula f(3) - f(0) all divided by 3

Vocaloid (vocaloid):

\[\frac{ f(3)-f(0) }{3 }\]

Vocaloid (vocaloid):

so, do you want to try calculating that for g(x) on your own?

OpenStudy (calculusxy):

I got: (3 - 0)/3 = 3/3 =1

OpenStudy (calculusxy):

I know. I did it in \(g(x) = x^2 - 2x\) I did \(g(0) = 0^2 - 2(0) = 0\) and \(g(3) = 3^2 - 2(3) = 3\)

Vocaloid (vocaloid):

good, now do the same thing for h and f

OpenStudy (calculusxy):

For \(f(x)\) it would be 6/3 = 2

Vocaloid (vocaloid):

good

Vocaloid (vocaloid):

h?

OpenStudy (calculusxy):

And for h(x) it would be 7/3 = 2.33333

Vocaloid (vocaloid):

good, so in order it's f, h and g

Vocaloid (vocaloid):

wait wait

OpenStudy (calculusxy):

There's no option like that

Vocaloid (vocaloid):

they want greatest to least

Vocaloid (vocaloid):

so reverse what i told you

Vocaloid (vocaloid):

g, h, then f

OpenStudy (calculusxy):

There's nothing of that

Vocaloid (vocaloid):

ah i got it

Vocaloid (vocaloid):

g = 1 f = 2 h = 2.3 so h, f ,g (4th option)

OpenStudy (calculusxy):

Thank you :)